By X-Ray diffraction method, the unit length of NaCl is observed to be 0.5627nm. The density of NaCl is found to be 2.164g/cm3. What type of defect exists in the Crystal? Calculate the percentage of Na+ and Cl- ions missing?

Grade:12

1 Answers

Supratim Das
6 years ago
DENSITY OF A UNIT CELL(D)= NO. OF ATOMS OF A UNIT CELL(Z) x ATOMIC MASS (M) AVOGADRO NUMBER (N) x VOLUME OF A UNIT CELL (a^3) D = 2.164 g/ccM = 23 + 35.5 = 58.5 ga = 0.5627 nm = 0.5627x 10^(-10) cmN = 6.023 x 10^(23)Z = 2.164 x 6.023 x 10^(23) x {0.5627x 10^(-10)}^(3) 58.5 = 3.96WE KNOW THAT NaCl HAS FCC STRUCTUREZ = 4THEREFORE, % OF MISSING IONS = 4 - 3.96 X 100 4 = 1%